Determine and Use Overhead Rate
Cortez Consulting, Inc., expects the following costs and expenses during the coming year:

Direct labor (@ $22.50/hr.) $81,000

Sales commissions

18,500
Overhead 72,000

a. Compute a predetermined overhead rate applied on the basis of direct labor hours.

Round to two decimal places, if applicable.

$Answer:

b. Prepare a general journal entry to apply overhead during an interim period when 750 direct labor hours were worked.

General Journal
Description Debit Credit
Answer Choices:Consulting overhead/Cost of goods sold/Finished goods inventory/Work in process inventory Answer Answer
Answer Choices:Consulting overhead/Cost of goods sold/Finished goods inventory/Work in process inventory Answer Answer

c. What amount of overhead would be assigned to Job 466, to which $90 in direct labor had been charged?

Round to two decimal places, if applicable.

$Answer:
Answer #1
(a)
Estimated Direct labor hours = Estimated Direct labor cost / Estimated Direct labor rate = 81000 / 22.50 3600 hours
Predetermined overhead rate = Estimated overhead / Estimated direct labor hours = 72000 / 3600 20.00
(b)
General Journal
Description Debit Credit
Work in process inventory ( 750*20 ) 15000
   Consulting overhead 15000
(c)
Direct labor hours worked on Job 466 = Direct labor cost charged to Job 466 / Direct labor rate = 90 / 22.50 4
Amount of overhead that would be assigned to Job 466 = Direct labor hours worked on Job 466 * Predetermined overhead rate = 4 * 20 80.00